100 ml of 0.3 N HCl solution is mixed with 200 ml of 0.6 N H2SO4 solution. What is the normality of H2SO4 in the solution?

Grade:12th pass

1 Answers

Vikas TU
6 years ago
From N1V1 = N2V2 we get,
100*0.3 + 200*0.6 = N*300
we get,
N = (30 + 120)/300 => 150/300 => 0.5
Thefore for H2SO4 the normality would become => 0.5*2 = 1 (2 being the valence factor)